# Development of ion-beam sputtered silicon nitride thin films for low-noise mirror coatings of gravitational-wave detectors

### v1

@@ 3.4.1 Density and phase
In order to estimate its density $\rho$, sample 22028 ($V=700$ V, $I=195$ mA, N2/Ar = 2) was used in XRR and RBS measurements. Fig. 10 shows the acquired XRR spectrum, together with its best-fit simulation where the sample was modeled as a thin SiNx layer on top of a Si substrate of semi-infinite thickness. The density value deduced from such simulation was 3.04 $\pm$ 0.01 g/cm3. Figure 5 shows the experimental and simulated RBS spectra. In this case, $\rho=3.038\pm 0.006$ g/cm3 was calculated using values of atom areal density and relative atomic concentrations, as deduced from stopping power simulations, and of thickness, as deduced from ellipsometric measurements.
@@
Sample 22028 was then treated in-air at 1000 C for 10 hours and used again in RBS and XRR measurements, to observe the effect of annealing. We found that annealing induced a slight increase of the sample density to $3.06\pm 0.01$ g/cm3, and at the same time an appreciable increase of the oxygen relative concentration, up to a value compatible with the formation of a 7-nm thick layer of stoichiometric SiO2 at the sample surface. As shown by the diffractogram presented on Fig. 11, the phase of the sample remained amorphous after annealing.
@@ 3.4.2 Coating loss angle
Figure 12 and Table 7 show the measured loss angle $\varphi$ of samples c21013, c21015 and d2008, produced using different sets of deposition parameters. Measurements were performed before and after in-air annealing at 900 C, and as a function of frequency. First we tested the effect of the ion beam composition, changing it from pure nitrogen to N2/Ar = 2, then we used two different values of ion beam current $I$ which yielded significantly lower optical absorption, $I=100$ and $I=195$ mA. For comparison, Fig. 12 also shows a data set of a sample deposited by low-pressure chemical-vapor deposition [^1] The following deposition parameters were used: NH3 mass flow of 60 sccm, SiH2Cl2 mass flow of 20 sccm, working pressure of 1.33 mbar, substrate temperature of 800 C..
@@
At approximately 2.8 kHz, the lowest resonant frequency of samples c21013 and c21015, we found that the loss angle of samples d2008 and c21013 after annealing is 2.2 and 3.3 times lower than that of tantala-titania layers in current mirror coatings of gravitational-wave interferometers [10], respectively, and appreciably lower than results found in the literature [22, 32]. The loss angle value measured on sample c21013 after annealing is the lowest ever measured on IBS SiNx thin films.
@@ 3.5 Mirror coatings
The use of multiple materials at different depths in a mirror coating, or multi-material design, is a design approach which allows to use thin layers of lower Brownian noise despite their higher optical absorption [66, 67]. We thus used our IBS SiNx thin films with lowest absorption and lower loss angle ($V=700$ V, $I=195$ mA, N2/Ar = 2) to realize a thickness-optimized ternary Bragg reflector, that is, a high-reflection mirror coating composed of stacked thin layers of three different materials. The other materials used were tantala-titania (higher loss angle, lower absorption) and silica (lower loss angle and absorption), as in current mirror coatings of gravitational-wave interferometers [10].
@@
Samples of our ternary mirror coating were deposited on fused silica substrates (25.4 mm diameter, 6 mm thickness), then annealed and measured for transmittance $T$, optical absorption at 1064 nm, and coating thermal noise. All resulted compliant with the target value of less than 5.6 ppm. By way of example, Table 9 lists the results of coating noise and absorption of sample 10039 as deposited and of sample 10036 after annealing. This latter features a noise amplitude at 100 Hz that is 25% lower than that of the current coatings of Advanced LIGO and Advanced Virgo, and only a slightly higher optical absorption.
@@ Acknowledgments
This work has been promoted by the LMA and partially supported by the Master Projet Couches Minces Optiques of CNRS Nucléaire & Particules, and by the European Gravitational Observatory in the framework of the Advanced Virgo+ project. The integrated XPS measurements were carried out at the SmartLab facility of the Department of Physics at Sapienza Università di Roma. We thank I. Roch and G. Cochez of the Institut d’Electronique, de Microélectronique et de Nanotechnologie for providing an LP-CVD sample, in the framework of the French RENATECH network. The CTN experiment is supported by the LIGO Laboratory. LIGO was constructed by the California Institute of Technology and Massachusetts Institute of Technology with funding from the National Science Foundation and operates under cooperative agreement PHY-1764464. CTN data analysis and figure generation were carried out using MATLAB software provided by The MathWorks, Inc. M. Magnozzi gratefully acknowledges the RADIATE Transnational Access grant n. 824096 (proposal 21002507-ST) for access to the SAFIR ion beam facility at the Institut des NanoSciences de Paris, funded by the Horizon 2020 research and innovation programme of the European Union.
In the document repositories of the LIGO and Virgo Collaborations, this work has numbers LIGO-P2400299 and VIR-0630B-24, respectively.

### v2

@@ 3.4.1 Density and phase
In order to estimate its density $\rho$, sample 22028 ($V=700$ V, $I=195$ mA, N2/Ar = 2) was measured using XRR and RBS. Figure 10 shows the acquired XRR spectrum, together with its best-fit simulation where the sample was modeled as a thin SiNx layer on top of a Si substrate of semi-infinite thickness. The density value deduced from such simulation was 3.04 $\pm$ 0.01 g/cm3. Figure 5 shows the experimental and simulated RBS spectra. In this case, $\rho=3.038\pm 0.006$ g/cm3 was calculated using values of atom areal density and relative atomic concentrations, as deduced from stopping power simulations, and of thickness, as deduced from ellipsometric measurements.
@@
Sample 22028 was then treated in-air at 1000 C for 10 hours and used again in RBS and XRR measurements, to observe the effect of annealing. We found that annealing induced a slight increase of the sample density to $3.06\pm 0.01$ g/cm3, and at the same time an appreciable increase of the oxygen relative concentration, up to a value compatible with the formation of a 7-nm thick layer of stoichiometric SiO2 at the sample surface. The full oxidation of the topmost layer at the sample surface was found to be in agreement with results found in the literature [65]. As shown by the diffractogram presented on Fig. 11, the phase of the sample remained amorphous after annealing.
@@ 3.4.2 Coating loss angle
Figure 12 and Table 7 show the measured loss angle $\varphi$ of samples c21013, c21015 and d2008, produced using different sets of deposition parameters. Measurements were performed before and after in-air annealing at 900 C, and as a function of frequency. First we tested the effect of the ion beam composition, changing it from pure nitrogen to N2/Ar = 2, then we used two different values of ion beam current $I$ which yielded significantly lower optical absorption, $I=100$ and $I=195$ mA (see Table 2). For comparison, Fig. 12 also shows a data set of a sample deposited by low-pressure chemical-vapor deposition (LP-CVD) [^1] For the LP-CVD sample, the following deposition parameters were used: NH3 mass flow of 60 sccm, SiH2Cl2 mass flow of 20 sccm, working pressure of 1.33 mbar, substrate temperature of 800 C..
@@
At approximately 2.8 kHz, which was the lowest resonant frequency accessible with our samples, we found that the loss angle of samples d2008 and c21013 after annealing is 2.2 and 3.3 times lower than that of tantala-titania layers in current mirror coatings of gravitational-wave interferometers [10], respectively, and appreciably lower than results found in the literature [22, 32]. To the best of our knowledge, the loss angle value measured on sample c21013 after annealing is the lowest ever measured on IBS SiNx thin films.
@@ 3.5 Mirror coatings
The use of multiple materials at different depths in a mirror coating, or multi-material design, is a design approach which allows to use thin layers of lower Brownian noise despite their higher optical absorption [68, 69]. We thus used our IBS SiNx thin films with lowest absorption and lower loss angle ($V=700$ V, $I=195$ mA, N2/Ar = 2) to design and realize a thickness-optimized ternary Bragg reflector, that is, a high-reflection mirror coating composed of stacked thin layers of three different materials [70]. The other materials used were tantala-titania (higher loss angle, lower absorption) and silica (lower loss angle and absorption), as in current mirror coatings of gravitational-wave interferometers [10].
@@
Samples of our ternary mirror coating were deposited on Corning 7980 fused silica substrates (25.4 mm diameter, 6 mm thickness), then annealed and measured for transmittance $T$, optical absorption at 1064 nm, and coating thermal noise. All resulted compliant with the target $T$ value of less than 5.6 ppm. By way of example, Table 9 lists the results of coating noise and absorption of sample 10039 as deposited and of sample 10036 after annealing. This latter features a noise amplitude at 100 Hz that is 25% lower than that of the current coatings of Advanced LIGO and Advanced Virgo. Such measured noise amplitude was found to be in agreement with the prediction of our design optimization code. Concerning the optical absorption, the average value of 1.6 ppm obtained by combining CTN and PTD results was also found to be in agreement with the prediction of our design optimization code.
Work is ongoing to characterize the scattering properties of our ternary mirror coating prototype.
@@ Acknowledgments
This work has been promoted by the LMA and partially supported by the Master Projet Couches Minces Optiques of CNRS Nucléaire & Particules and by the European Gravitational Observatory in the framework of the Advanced Virgo+ project.
The CTN experiment is supported by the LIGO Laboratory. LIGO was constructed by the California Institute of Technology and Massachusetts Institute of Technology with funding from the National Science Foundation and operates under cooperative agreement PHY-1764464. CTN data analysis and figure generation were carried out using MATLAB software provided by The MathWorks, Inc.
M. Magnozzi gratefully acknowledges the RADIATE Transnational Access grant n. 824096 (proposal 21002507-ST) funded by the Horizon 2020 research and innovation programme of the European Union, for access to the SAFIR ion beam facility at the Institut des NanoSciences de Paris.
The integrated XPS measurements were carried out at the SmartLab facility of the Department of Physics at Sapienza Università di Roma.
We thank I. Roch and G. Cochez of the Institut d’Electronique, de Microélectronique et de Nanotechnologie for providing an LP-CVD sample, in the framework of the French RENATECH network.
In the document repositories of the LIGO and Virgo Collaborations, this work has numbers LIGO-P2400299 and VIR-0630C-24, respectively.
Data availability
The data that support the findings of this article are available from the authors upon reasonable request.